OpenStudy (anonymous):

What is the equation of the line that passes through (–2, 3) and is parallel to 2x + 3y = 6?

OpenStudy (anonymous):

If you know that the equation is parallel to 2x+3y=6, it must have the same slope (so that it'll never touch, right?). To find the slope of the given line, put it into slope-intercept form, so you get 3y=-2x+6 y=-2/3x+2 slope= -2/3 substitute that into y=mx+b, as the b-value. y=-2/3x+b now to solve for the equation that runs through the point (-2, 3), you need to substitute that as the x- and y-values and solve for b 3=(-2/3)(-2)+b 3=4/3+b 5/3=b substitute the b-value into the equation where it is just y=-2/3x+b to get your final function y=-2/3x+5/3 and if you need to put this in standard form, it becomes y+2/3x=5/3 3y+2x=5

OpenStudy (anonymous):

but these are the answers

OpenStudy (anonymous):

A.y=-3/2x -6 B.y=-2/3x + 6 C.y=2x + 3 D.y=3/2x